OpenStudy (anonymous):

expressed in simplest radical form the, the product of radical 6 * radical 15 is

OpenStudy (mertsj):

\[\sqrt{6}\sqrt{15}\sqrt{6\times15}=\sqrt{2\times3^2\times5}=3\sqrt{10}\]